# BEP: A Binary Error Propagation Algorithm for Binary Neural Networks Training

### v1

@@ 4.1 Experimental Setup
Our evaluation spans a diverse set of benchmarks. For the feedforward MLPs, we adopt the same datasets as (colombo2025training): Random Prototypes, FashionMNIST (xiao2017fashion), CIFAR-10 (krizhevsky2009learning), and Imagenette (imagenette). The Random Prototypes dataset comprises 20,000 training and 3,000 test samples with input dimension $K_{0}=1000$ and a flip probability $p=0.46$. FashionMNIST contains 50,000 training and 10,000 test grayscale images of size $28\times 28$. For CIFAR-10 and Imagenette, we use the same AlexNet-derived features as in (colombo2025training). Both are 10-class datasets; CIFAR-10 has 50,000 training and 10,000 test samples, while Imagenette has 10,000 training and 1,963 test samples. To evaluate recurrent models, we use Sequential MNIST (S-MNIST) (deng2012mnist) and 30 real-world time-series classification tasks from the UCR Archive (dau2019ucr), which cover a wide range of sample sizes, feature dimensions, and class counts. In all experiments, we utilize a fixed output classifier $\mathbf{P}$ generated via the binary equiangular frame method detailed in Appendix C.
@@
Figure 2: Test accuracy as a function of the number of parameters on Random Prototypes, FashionMNIST, CIFAR10, and Imagenette. Results compare BEP with both the SotA approach (colombo2025training) and QAT-based methods for binary MLPs with $L=2$ and $L=3$ hidden layers.
@@ 4.2 Performance on Binary Multi-Layer Perceptrons
Our first experiment evaluates BEP against the current SotA approach for binary training of MLPs (colombo2025training). Specifically, we considered binary MLPs with two and three hidden layers, trained using both BEP and the SotA local learning rule. The tuned hyperparameters, across $e=50$ epochs, include the robustness parameter $r$, reinforcement probability $p_{r}$, initial group size $\gamma_{0,l}$, and gating threshold $\nu$. For a comprehensive benchmark, we also trained the same architectures using the QAT implementation from the Larq framework (geiger2020larq). In this setup, the forward pass employs binarized weights, while the backward pass updates latent FP parameters with the Adam optimizer (kingma2014adam). To ensure a fair comparison, the main QAT baseline does not use batch normalization, resulting in a fully binary model at inference. For completeness, we also report the performance of QAT with batch normalization as a reference point.
@@
Table 1: Test accuracy on 30 UCR datasets. The results compare our BEP algorithm with the QAT-based training from the Larq framework (geiger2020larq) for binary RNN architectures.
@@ 4.3 Validation on Binary Recurrent Neural Networks
Our second experiment tests the BEP algorithm on binary RNNs using its time-unrolled variant BEP-TT described in Section 3.5. The goal is to demonstrate that a global, end-to-end error signal can effectively train recurrent models, a task typically intractable for purely local learning rules. Specifically, we evaluate BEP-TT on many-to-one sequence classification tasks across 30 datasets from the UCR Time Series Archive (dau2019ucr), using only the last window of each time series. All RNN results use 3-fold cross-validation and 3 independent runs with the same hyperparameters: robustness $r=0.5$, reinforcement probability $p_{r}=0.5$, initial group size $\gamma_{0,l}=15$, gating threshold $\nu=0.05$, hidden and output layer sizes $K_{s}=K_{y}=1035$, training epochs $e=50$, and batch size $bs=N/10$. As a baseline, we trained binary RNNs with the same architecture using the QAT implementation from the Larq framework (geiger2020larq). As in the previous experiment, no batch or layer normalization was applied so that the models remain fully binary at inference time. For completeness, we report the performance of QAT with batch normalization as a reference point. For all datasets and models, inputs were binarized with a distributive thermometer encoder (pmlr-v235-bacellar24a), followed by a fixed $\pm 1$ expansion layer projecting the dimension to $K_{0}=1035$. A hyperparameter search was conducted to tune the sequence window length, defined as the number of timesteps included in the temporal window, and the number of thermometer bits.
@@ 4.5 Discussion and Limitations
A key advantage of BEP lies in its ability to perform training using only bitwise computations. Table 2 summarizes the memory footprint and computational complexity of BEP compared with standard QAT. Relative to QAT-based methods, our approach achieves a $2\times$ reduction in memory usage for hidden weights and a $32\times$ reduction for error signals and weight updates. Moreover, Adam-based QAT requires an additional $64$ bits per parameter to store its first- and second-order moment estimates, further increasing its memory demand. From a computational standpoint, we adopt the hardware-level metric of equivalent Boolean gates (colombo2025training) to compare the intrinsic cost of arithmetic operations. All operations in BEP can be implemented using primitives such as XNOR, Popcount, and increment/decrement operations, which require at most $\mathcal{O}(10N)$ Boolean gates for $N$-bit operands (bacellar2024differentiable). In contrast, IEEE-754 single-precision FP additions and multiplications (8766229), which underpin Adam-based QAT, are estimated to require on the order of $10^{4}$ Boolean gates each (luo2024addition). This comparison highlights the substantial computational advantage of BEP, which reduces the Boolean-gate cost by approximately three orders of magnitude relative to FP32 QAT with Adam.
Despite its promising results, our approach has some limitations that suggest natural directions for future research. First, we focus primarily on binary MLPs and RNNs. Extending BEP to convolutional or transformer-style models requires a full binary design for these architectures, including handling binary convolutions, filter-level masking, and additional adaptations to support weight sharing, spatial structure, and multi-head mechanisms. Second, our experiments are restricted to classification tasks. Although BEP naturally supports arbitrary binary output vectors and could, in principle, be applied to multi-label prediction, binary segmentation, or more general binary-vector regression, such tasks necessitate additional design choices regarding the output encoding. Third, we evaluate BEP on mid-scale datasets and moderate-depth NNs. Extending BEP to large-scale models (e.g., ImageNet-level CNNs (bulat2019xnornetpp; sarkar2024biper)) requires substantial architectural adaptations to convolutional or transformer backbones. Moreover, binary-compatible normalization mechanisms may become necessary to ensure stable training in such settings.
Table 2: Comparison of memory footprint and computational complexity between QAT (with Adam) and BEP. Memory is reported in bits per element, and complexity denotes the order of magnitude of equivalent Boolean-gate operations per element (colombo2025training). : This estimate assumes the absence of batch normalization and full-precision scaling factors commonly used in QAT.
@@ 5 Conclusion and Future Work
In this paper, we introduced BEP, an algorithm for training multi-layer BNNs using exclusively binary computations. The central contribution is the formulation of a principled, binary analog of the BP algorithm. By defining a recursive rule for propagating binary-valued error signals and updating integer-valued metaplastic weights, BEP bridges the gap between the global credit assignment of gradient-based learning and the computational efficiency of bitwise operations. Theoretically, this work shows that effective end-to-end learning in multi-layer binary NNs is possible without relying on continuous gradients, opening new avenues for analyzing discrete optimization in DL. Practically, BEP enables efficient training using only XNOR, Popcount, and increment/decrement operations, making it well-suited for constrained settings such as TinyML (capogrosso2024machine; pavan2024tybox), privacy-preserving DL with homomorphic encryption (falcetta2022privacy; colombo2024training), and neuromorphic systems (indiveri2015memory; yamazaki2022spiking). Future work includes extending BEP to convolutional architectures, refining adaptive strategies for the gating threshold $\nu$, exploring learnable masking mechanisms, and developing a formal convergence analysis.
@@ Appendix C Generating a Fixed Binary Classifier via Equiangular Frames
As stated in Section 3, our empirical results show that BEP achieves its best performance when using a fixed output classifier $\mathbf{P}$ whose class prototypes are geometrically well-separated. This approach is inspired by the concept of Equiangular Tight Frames (ETFs), which have been shown to emerge in the final layers of deep NNs during a phenomenon known as neural collapse (Papyan2020). While a simple, randomly generated classifier offers a baseline, optimizing the structure of these prototypes significantly improves class separability. This appendix details our method for generating a structured binary classifier by constructing a set of prototype vectors that are maximally and uniformly distant from each other in the binary feature space.

### v2

@@ 4.1 Experimental Setup
Our evaluation spans a diverse set of benchmarks. For the feedforward MLPs, we adopt the same datasets as (Colombo et al., 2025): Random Prototypes, FashionMNIST (Xiao et al., 2017), CIFAR-10 (Krizhevsky et al., 2009), and Imagenette (Howard, 2019). The Random Prototypes dataset comprises 20,000 training and 3,000 test samples with input dimension $K_{0}=1000$ and a flip probability $p=0.46$. FashionMNIST contains 50,000 training and 10,000 test grayscale images of size $28\times 28$. For CIFAR-10 and Imagenette, we use the same AlexNet-derived features as in (Colombo et al., 2025). Both are 10-class datasets; CIFAR-10 has 50,000 training and 10,000 test samples, while Imagenette has 10,000 training and 1,963 test samples. To evaluate recurrent models, we use Sequential MNIST (S-MNIST) (Deng, 2012) and 30 real-world time-series classification tasks from the UCR Archive (Dau et al., 2019), which cover a wide range of sample sizes, feature dimensions, and class counts. In all experiments, we utilize a fixed output classifier $\mathbf{P}$ generated via the binary equiangular frame method detailed in Appendix C.
@@
Figure 2: Test accuracy as a function of the number of parameters on Random Prototypes, FashionMNIST, CIFAR10, and Imagenette. Results compare BEP with both the SotA approach (Colombo et al., 2025) and QAT-based methods for binary MLPs with $L=2$ and $L=3$ hidden layers.
@@ 4.2 Performance on Binary Multi-Layer Perceptrons
Our first experiment evaluates BEP against the current SotA approach for binary training of MLPs (Colombo et al., 2025). Specifically, we considered binary MLPs with two and three hidden layers, trained using both BEP and the SotA local learning rule. The tuned hyperparameters, across $e=50$ epochs, include the robustness parameter $r$, reinforcement probability $p_{r}$, initial group size $\gamma_{0,l}$, and gating threshold $\nu$. For a comprehensive benchmark, we also trained the same architectures using the QAT implementation from the Larq framework (Geiger and Team, 2020). In this setup, the forward pass employs binarized weights, while the backward pass updates latent FP parameters with the Adam optimizer (Kingma and Ba, 2014). To ensure a fair comparison, the main QAT baseline does not use batch normalization, resulting in a fully binary model at inference. For completeness, we also report the performance of QAT with batch normalization as a reference point.
@@
Table 1: Test accuracy on 30 UCR datasets. The results compare our BEP algorithm with the QAT-based training from the Larq framework (Geiger and Team, 2020) for binary RNN architectures.
@@ 4.3 Validation on Binary Recurrent Neural Networks
Our second experiment tests the BEP algorithm on binary RNNs using its time-unrolled variant BEP-TT described in Section 3.5. The goal is to demonstrate that a global, end-to-end error signal can effectively train recurrent models, a task typically intractable for purely local learning rules. Specifically, we evaluate BEP-TT on many-to-one sequence classification tasks across 30 datasets from the UCR Time Series Archive (Dau et al., 2019), using only the last window of each time series. All RNN results use 3-fold cross-validation and 3 independent runs with the same hyperparameters: robustness $r=0.5$, reinforcement probability $p_{r}=0.5$, initial group size $\gamma_{0,l}=15$, gating threshold $\nu=0.05$, hidden and output layer sizes $K_{s}=K_{y}=1035$, training epochs $e=50$, and batch size $bs=N/10$. As a baseline, we trained binary RNNs with the same architecture using the QAT implementation from the Larq framework (Geiger and Team, 2020). As in the previous experiment, no batch or layer normalization was applied so that the models remain fully binary at inference time. For completeness, we report the performance of QAT with batch normalization as a reference point. For all datasets and models, inputs were binarized with a distributive thermometer encoder (Bacellar et al., 2024a), followed by a fixed $\pm 1$ expansion layer projecting the dimension to $K_{0}=1035$. A hyperparameter search was conducted to tune the sequence window length, defined as the number of timesteps included in the temporal window, and the number of thermometer bits.
@@ 4.5 Discussion and Limitations
A key advantage of BEP lies in its ability to perform training using only bitwise computations. Table 2 summarizes the memory footprint and computational complexity of BEP compared with standard QAT. Relative to QAT-based methods, our approach achieves a $2\times$ reduction in memory usage for hidden weights and a $32\times$ reduction for error signals and weight updates. Moreover, Adam-based QAT requires an additional $64$ bits per parameter to store its first- and second-order moment estimates, further increasing its memory demand. From a computational standpoint, we adopt the hardware-level metric of equivalent Boolean gates (Colombo et al., 2025) to compare the intrinsic cost of arithmetic operations. All operations in BEP can be implemented using primitives such as XNOR, Popcount, and increment/decrement operations, which require at most $\mathcal{O}(10N)$ Boolean gates for $N$-bit operands (Bacellar et al., 2024b). In contrast, IEEE-754 single-precision FP additions and multiplications (21), which underpin Adam-based QAT, are estimated to require on the order of $10^{4}$ Boolean gates each (Luo and Sun, 2024). This comparison highlights the substantial computational advantage of BEP, which reduces the Boolean-gate cost by approximately three orders of magnitude relative to FP32 QAT with Adam.
Despite its promising results, our approach has some limitations that suggest natural directions for future research. First, we focus primarily on binary MLPs and RNNs. Extending BEP to convolutional or transformer-style models requires a full binary design for these architectures, including handling binary convolutions, filter-level masking, and additional adaptations to support weight sharing, spatial structure, and multi-head mechanisms. Second, our experiments are restricted to classification tasks. Although BEP naturally supports arbitrary binary output vectors and could, in principle, be applied to multi-label prediction, binary segmentation, or more general binary-vector regression, such tasks necessitate additional design choices regarding the output encoding. Third, we evaluate BEP on mid-scale datasets and moderate-depth NNs. Extending BEP to large-scale models (e.g., ImageNet-level CNNs (Bulat and Tzimiropoulos, 2019; Vargas et al., 2024)) requires substantial architectural adaptations to convolutional or transformer backbones. Moreover, binary-compatible normalization mechanisms may become necessary to ensure stable training in such settings.
Table 2: Comparison of memory footprint and computational complexity between QAT (with Adam) and BEP. Memory is reported in bits per element, and complexity denotes the order of magnitude of equivalent Boolean-gate operations per element (Colombo et al., 2025). : This estimate assumes the absence of batch normalization and full-precision scaling factors commonly used in QAT.
@@ 5 Conclusion and Future Work
In this paper, we introduced BEP, an algorithm for training multi-layer BNNs using exclusively binary computations. The central contribution is the formulation of a principled, binary analog of the BP algorithm. By defining a recursive rule for propagating binary-valued error signals and updating integer-valued metaplastic weights, BEP bridges the gap between the global credit assignment of gradient-based learning and the computational efficiency of bitwise operations. Theoretically, this work shows that effective end-to-end learning in multi-layer binary NNs is possible without relying on continuous gradients, opening new avenues for analyzing discrete optimization in DL. Practically, BEP enables efficient training using only XNOR, Popcount, and increment/decrement operations, making it well-suited for constrained settings such as TinyML (Capogrosso et al., 2024; Pavan et al., 2024), privacy-preserving DL with homomorphic encryption (Falcetta and Roveri, 2022; Colombo et al., 2024), and neuromorphic systems (Indiveri and Liu, 2015; Yamazaki et al., 2022). Future work includes extending BEP to convolutional architectures, refining adaptive strategies for the gating threshold $\nu$, exploring learnable masking mechanisms, and developing a formal convergence analysis.
Acknowledgements
This paper is supported by Dhiria S.r.l. and by PNRR-PE-AI FAIR project funded by the NextGeneration EU program.
@@ Appendix C Generating a Fixed Binary Classifier via Equiangular Frames
As stated in Section 3, our empirical results show that BEP achieves its best performance when using a fixed output classifier $\mathbf{P}$ whose class prototypes are geometrically well-separated. This approach is inspired by the concept of Equiangular Tight Frames (ETFs), which have been shown to emerge in the final layers of deep NNs during a phenomenon known as neural collapse (Papyan et al., 2020). While a simple, randomly generated classifier offers a baseline, optimizing the structure of these prototypes significantly improves class separability. This appendix details our method for generating a structured binary classifier by constructing a set of prototype vectors that are maximally and uniformly distant from each other in the binary feature space.